X is 6 years younger than Y, and the average of their ages is twice X age 5 years ago, how old are they?
Answer by LilSkittleMd(119)

First set up equations
x=y-6
(x+y)/2=2x-5
Substitute "y-6" where any x's are
(y-6+y)/2=2(y-6)-5
Combine like terms,distribute
2y-6/2=2y-12-5
Factor 2y-6 to cancel the 2 so there is no fraction
2(y-3)/2=2y-17
The two 2's cancel each other out so you are left with
y-3=2y-17
subtract y from both sides
-3=y-17
add 17 to both sides
14=y
x=14-6
x=8
X is 8 years old, Y is 14 years old